OpenStudy (anonymous):

Find the area between the graph of y=x^2-2 and the x-axis, between x=0 and x=5. I took the integral and got 31.667.. not sure if thats right

OpenStudy (dumbcow):

\[\int\limits_0^5 x^2 - 2 = \frac{x^3}{3} - 2x \rightarrow \frac{5^3}{3} - 10\]

OpenStudy (dumbcow):

you are correct

OpenStudy (anonymous):

thanks!